OpenStudy (anonymous):

x+{y-3-4x+2[3x-(-y-1)+2y]-3x} = ?

OpenStudy (anonymous):

i am having a little trouble solving this one

OpenStudy (turingtest):

one thing is that it's a lot easier on the eyes if you don't use caps

OpenStudy (anonymous):

wow can you explain how to do that

OpenStudy (anonymous):

x+{y-3-4x+2[3x-(-y-1)+2y]-3x} = ?

OpenStudy (turingtest):

I edited your post for you :) anyway, we work here from the inside out so what is the first (innermost) operation?

OpenStudy (anonymous):

i saw that

OpenStudy (anonymous):

(-y-1)

OpenStudy (turingtest):

right, and there's not much we can do with that... so on to the next step we now look at it as\[3x-(-y-1)+2y\]do you know how to simplify this?

OpenStudy (anonymous):

i believe it is 3x+y+1+2y then 3x+2y+1

OpenStudy (turingtest):

yep :)

OpenStudy (anonymous):

i meant 3y

OpenStudy (turingtest):

ph, so did I lol

OpenStudy (turingtest):

so what is the next step outside of that?

OpenStudy (anonymous):

3x+3y+1

OpenStudy (turingtest):

x+{y-3-4x+2[3x-(-y-1)+2y]-3x} and we got that now to be x+{y-3-4x+2[3x+3y+1]-3x} so what is the innermost step we should do (careful on this one)

OpenStudy (anonymous):

{y-3-4x+2[3x+3y+1]-3x}

OpenStudy (anonymous):

right?

OpenStudy (turingtest):

almost.... and that can work, but we can make it even simpler the first thing to be done (the innermost operation) is to distribute the 2: 2[3x+3y+1]

OpenStudy (anonymous):

yah i know thats what i meant

OpenStudy (turingtest):

good so simplify that and tell me what you get

OpenStudy (anonymous):

ok we will see?

OpenStudy (anonymous):

2[3x+3y+1] which is 6x +6y +2

OpenStudy (turingtest):

good, so now substitute that into the next outer expression (the one you wrote before)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x+y-3-4x+6x +6y +2-3x

OpenStudy (turingtest):

great, and now we simplify that....

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

+7y - 1?

OpenStudy (anonymous):

am i right?

OpenStudy (turingtest):

no be more careful with your like terms what are just all the x's added up?

OpenStudy (turingtest):

oh wait I'm sorry I mistyped my own paper (mulitptasking)

OpenStudy (anonymous):

:( my book says i am right

OpenStudy (turingtest):

oh I see you took the last step already which is to add the final outer x well good job, yes, that is the final answer

OpenStudy (anonymous):

ok well can we try one more to make sure i learned everything well. if not thats fine you earned your medal?

OpenStudy (turingtest):

I'm never in it for the medals :) but you're still supposed to post separate each Q as per the code of conduct In this case we can do one more quickly here as I said, I'm multitasking

OpenStudy (anonymous):

ok lol i guess i wouldnt be in the medals either if i had your score

OpenStudy (anonymous):

ok next one

OpenStudy (turingtest):

go ahead

OpenStudy (anonymous):

6+{-a-3[b+6(a-b)+4a]-2b=?

OpenStudy (turingtest):

where to the brackets close {} ? (you forgot the right one)

OpenStudy (turingtest):

where do*....

OpenStudy (anonymous):

oh my bad after 2b

OpenStudy (anonymous):

6+{-a-3[b+6(a-b)+4a]-2b}=?

OpenStudy (turingtest):

so what is the first innermost operation?

OpenStudy (anonymous):

a-b

OpenStudy (turingtest):

right, but we can't simplify that, so what's the next outer operation (again, like last time, be careful here)

OpenStudy (anonymous):

6+{-a-3[b+6(a-b)+4a]-2b}=? then [b+6a-6b+4a]

OpenStudy (turingtest):

yes

OpenStudy (turingtest):

continue

OpenStudy (anonymous):

ok so now 6+{-a-3[b+6a-6b+4a]-2b}=? then [10a-5b] then -3[10a-5b]= -30a+15b

OpenStudy (turingtest):

so far so good

OpenStudy (anonymous):

so now we have 6+{-31a+17b}=? solve it now 6-31a+17b fix it up a bit... -31a+17b+6 am I right ,am I ,am I?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

6+{-31a+13b}=? solve it now 6-31a+13b fix it up a bit... -31a+13b+6 am I right ,am I ,am I?

OpenStudy (anonymous):

check that out

OpenStudy (turingtest):

yep, that's right :)

OpenStudy (anonymous):

alright thank you very much for your time